Determine whether ~u and ~v are orthogonal vectors, make an acute or obtuse angle: (1.1) ~u =< 1, 3, −2 >, ~v =< −5, 3, 2 >. (2) (1.2) ~u =< 1, −2, 4 >, ~v =< 5, 3, 7 >.


1
Expert's answer
2021-07-20T17:22:17-0400

1.1

"u=(1,3,-2), v=(-5,3,2)\\
uv=-5+9-4=0\\"

u orthogonal v


1.2

"u=(1,-2,4), v=(5,3,7)\\
uv=5-6+28=27>0"

u is not orthogonal v

"||u||=\sqrt{1+4+16}=\sqrt{21}\\
||v||=\sqrt{25+9+49}=\sqrt{83}\\
\cos\phi=\frac{uv}{||u||\cdot||v||}=\frac{27}{\sqrt{21}\cdot\sqrt{83}}>0"

acute angle